# LAVA: Explainability for Unsupervised Latent Embeddings

### v1

@@ A.1.3 Optimizing locality placement
Additionally, the DIRECT algorithm requires a range of values for the parameters in which the loss function is convex. We fix this to [4,4][-4,4] for both $\alpha$ and $\beta$. We conjecture that, at least approximately (especially since $\ell$-means is stochastic), the loss function is convex in that interval and contains a good solution. Loosely, as mentioned above, we want the weighing scheme to be able to overpower the overrepresentation of low-density regions that might occur if we would simply weigh samples based on their centrality. We argue about it in terms of hyperball volume: the hyperball volume is proportional to the $D_{L}$-th power of its radius, which for a neighborhood of a sample we can approximate by the average distance to its neighbors, where $D_{L}$ is the dimensionality of the latent space. To battle overrepresentation, we then want to weigh samples with larges “volumes” (samples in low density regions) inversely to their volume, so as to effectively force $\ell$-means to give them less importance than they might otherwise have given the squared Euclidean distance optimization. In the two-dimensional case, the hyperball volume is simply the area of the circle, which is proportional to the square of the radius (the average distance of the $n$ nearest neighbors). To allow weighing in such a fashion, we allow $\alpha$ to be at least $2$, and we further multiply by $2$ to allow for exaggeration of this principle. Since we cannot be sure how this weighing will interact with centrality weighing, we allow both parameters to also be negative, and give them the same range, arbitrarily. In our limited experimentation, and with no ground truth to compare to, this seemed to yield good and intuitive results (Figures S5 and S11). If we were to use a higher-dimensional latent space, we would set the range to $2D_{L},2D_{L}]$.
@@
The pseudocode, ommiting details about DIRECT, is shown in Algorithm S1. We use the scipy [^1] https://scipy.org/. implementation of the DIRECT algorithm, and the scikit-learn [^2] https://scikit-learn.org/. implementation of $k$-means.
$\mathbf{X}\in\mathbb{R}^{E\times D_{L}}$, $\ell$
@@
Algorithm S1 Locality placement algorithm.
@@ A.2 Locality Representation
Each locality’s correlations are calculated using the standard formulation of the Spearman correlation:
$r_{s}(\mathbf{X}_{i},\mathbf{X}_{j})=\frac{cov\left[\text{R}(\mathbf{X}_{i}),\text{R}(\mathbf{X}_{j})\right]}{\sigma_{\text{R}(\mathbf{X}_{i})},\sigma_{\text{R}(\mathbf{X}_{j})}},$ (9)
where $X_{i}$ and $X_{j}$ are two original input features and R is a ranking function. Covariance covcov and standard deviations $\sigma$ are estimated from the samples within a locality using standard estimators. Overlapping ranks are averaged.
Importantly, if one of the variables has the same value in $75\%$ of the neighborhood, that correlation is set to 0 for that locality. This is done to prevent overestimation of correlations in cases where there is very little change. Notably, this has been the case for both the MNIST and KPMP datasets in our experiments, as the values of most pixels in MNIST are 0, while gene expression at the single-cell level in KPMP is generally also very sparse. The value of $75\%$ was set arbitrarily, to allow for some leniency, but also avoid meaningless correlations based on low numbers of samples.
@@ A.3.1 AMF loss function and hyperparameters
Different metrics of (dis)similarity. We decided to use cosine distance for the main term of our loss function. However, a different metric (or divergence) for assessing multivariate (dis)similarity might be equally suitable, e.g., the Kullback-Leibler divergence. Pending further investigation, we did not assess whether the semantics of one of these might be more appropriate for our task than the other. Some work relating to semantics and efficiency of these metrics suggests triangular distance as a more appropriate choice in most contexts [51].
Up-weighing overestimation and connection to quantile regression. The $\nu$ hyperparameter in Equation 6 determines how up-weighted overestimation errors will be in our loss. Weighing the errors is done similarly to how it is done in univariate quantile regression. The loss for univariate $\tau$-th quantile regression, also known as the pinball loss [52], is defined as:
$\mathcal{L}_{\tau}(y,f(x))=\begin{cases}\tau\,|y-f(x)|&\text{if }y\geq f(x)\\ (1-\tau)\,|y-f(x)|&\text{if }y<f(x)\end{cases}.$ (10)
Practically speaking, univariate quantile regression produces a weighing scheme between overestimation and underestimation errors: if $\tau>0.5$, we are less concerned with overestimating $y$ using $f(x)$, and vice versa. Using a $\tau=0.5$ is equal to doing median regression, which corresponds to a regression that minimizes the mean absolute error rather than using the more standard mean squared error. For more intuition, we recommend reading [37] and [52].
The idea for module extraction is to retain the semantics of similar, scale-invariant patterns we get with cosine distance or a similar metric, but allowing for reasoning about the reconstructed correlations in terms similar to quantile regression. Retaining the error weighing with a $\tau<0.5$ is therefore done to prevent averaging across samples and extract only shared, reoccurring correlations (Section 2.3). Quantile-regression-like reconstructions will allow us to say that correlations present in module-based reconstructions are actually present in the localities with some certainty, or, in terms of errors, that the value of a correlation will be overestimated by at most $\tau$ (approximately). We can quantify whether this is actually the case by looking at the overestimation errors (Figure 4), which we calculate as the amount of overestimation errors divided by the total absolute error.
@@
$\centering\nu=\frac{1-\tau}{\tau}.\@add@centering$ (11)
@@
Using $\nu$ for weighing overestimation errors when using the mean absolute errors would be equivalent to doing $\tau$-th quantile regression. We have set the hyperparameter $\nu=9$ in all of our experiments. If we look at the pinball equivalent $\tau$, this would correspond to weighing in $0.1$-th quantile regression. We cannot make direct parallels since we do not optimize the mean absolute error, but we can observe relative amounts of overestimation, which seem to more-or-less correspond to values we would expect if we would to quantile regression (Figure 4).
@@
Extra: Presence fine-tuning. While we do not use it in our experiments, our implementation also includes what we call presence fine-tuning. After selecting modules, we can fix the module matrix $\mathbf{M}$ and optimize the presence matrix $\mathbf{P}$ with a pinball loss function (Equation 10), to guarantee that each locality (independent when $\mathbf{M}$ is fixed) has an overestimation ratio exactly equal to some $\tau$. Notice that the modules we find are contingent on the overestimation weighing used in the main loss term, which means that it is not equivalent to first extract, and then simply do fine-tuning at the end.
@@
Scale regularization. The second term in Equation 6 is a simple mean absolute error term designed to retain scale of reconstruction. The idea behind scaling is simply to make reconstructions of the same scale as the actual locality correlations, as cosine distance is in itself scale invariant. While the cosine similarity should remain relatively unaffected by this regularization, given a low enough hyperparameter $\gamma$, matching scale of the actual locality correlations and our reconstructions will allow us to observe the overestimation ratio of solutions (Section 3; Figure 4). Furthermore, due to scale invariance of cosine distance, it is also imaginable that modules avoid overestimation simply by retaining low correlations values across the board: the scaling should push optimization away from such solutions. It is crucial that the scale regularization contains no overestimation weighing, meaning that the ratio of overestimation is purely a result of the up-weighing in the first term of the loss function. This ensures the retainment of the semantics of reoccurring pattern semantics, rather than focusing on absolute values of our solutions. We set $\gamma=0.0001$ as the default value for scale regularization to have a low impact on the loss function. We also end up manually setting this hyperparameter for the KPMP dataset in the experiments (see Section B.3.2), as we noticed $\gamma=0.001$ had no effect on optimization because of the low amounts of observed correlation.
@@
Optimization hyperparameters. The matrices $\mathbf{P}$ and $\mathbf{M}$ are optimized to minimize the loss function (Equation 6) using gradient descent with minibatches of $64$ samples (arbitrarily set), the ADAM optimizer [53] with default hyperparameters, and clamping. Clamping is used to force any entries of matrices $\mathbf{P}$ and $\mathbf{M}$ to be within the $[0,1]$ interval after each step in training. Training is performed until convergence, which we defined as no improvement in the loss function above $1\%$ in $100$ epochs. After convergence, we use the matrices from the iteration with the lowest loss as the final solutions. We developed the code based on the PyTorch library [^4] https://pytorch.org/..
@@ B.1 Software and hardware
All our code was developed using Python 3.8 and its libraries [^5] Code included with supplementary material and will be made available online, conditional on acceptance.. We ran experiments on Linux compute nodes, comprising of a number of different CPUs and GPUs, details of which we omit to ensure anonymity. Locality placement and correlation calculations took several hours for both datasets. A single module extraction run would take about $5$ to $30$ minutes, increasing with the number of modules, $M$. We also tested all of our code on an Apple M3 Pro with 12 CPU cores and 18GB of RAM, which is what we also used to run Jupyter notebooks for result analysis and plot generation.
@@ B.2.1 MNIST
The MNIST dataset was downloaded using the Scikit-learn Python library [^6] https://scikit-learn.org. library. No preprocessing was performed on the data. Prior to running LAVA, we created a UMAP embedding of the dataset using the UMAP [^7] https://umap-learn.readthedocs.io. library, with default hyperparameters.
@@
The KPMP single-cell kidney atlas was downloaded from the CELLxGENE website [^8] https://cellxgene.cziscience.com. Specific dataset version: https://datasets.cellxgene.cziscience.com/f5b6d620-76df-45c5-9524-e5631be0e44a.h5ad.. The dataset contains a UMAP embedding of the data preprocessed using the Seurat library [^9] https://satijalab.org/seurat/., which we used for our experiments. More information can be found at the KPMP data portal website [^10] https://atlas.kpmp.org/.. The UMAP of the data can also be explored on the CELLxGENE website [^11] https://cellxgene.cziscience.com/e/dea717d4-7bc0-4e46-950f-fd7e1cc8df7d.cxg/.. For module extraction, we removed mitochondiral, ribosomal, and mitochondrial ribosome genes. Then, we selected the top $1000$ most variable genes, based on the gene expression data (after normalization and log transformation).
@@ B.3.1 Size of locality, $n$
Data-driven approach (MNIST). A data-driven way of choosing the value $n$ can be looking at the Jaccard similarity of neighborhoods of varying sizes in the original space and the latent space. Even though methods such as UMAP have hyperparameters that represent the size of neighborhoods that will be optimized for, different neighborhood sizes might end up being optimized (depending on the choice of other hyperparameters and the dataset). For the UMAP of the MNIST dataset used in our experiments (Section 3), we found that with the default hyperparameter of $15$ nearest neighbors, the resulting embedding neighborhoods were most representative of the neighborhoods in the original feature space when $k\approx 3000$ (Figure S1), which was the value we used for LAVA’s hyperparameter in our experiments with MNIST.
@@ B.3.2 Module extraction hyperparameters
For module extraction on both MNIST and KPMP, we use default hyperparameters, as defined in Appendix A.3, with one exception: the $\alpha$ regularization hyperparameter from Equation 6 in KPMP module extraction. $\boldsymbol{\alpha}$ for KPMP module extraction. In our runs, we noticed that using low values for $\alpha$ resulted in high overestimation error ratios. This pointed to the fact that regularization was mostly likely not affecting the loss function. We assumed this was due to the sparsity of gene expression, and subsequent low correlations in many localities. Running module extraction several times, using higher values of $\alpha$, confirmed this hypothesis (Figure S2). Increasing the strength of regularization seemed to have no or negative effect on the quality or stability of solutions (the main loss term and silhouette width, respectively). For our experiments, we ended up using $\alpha=10$, resulting in relatively stable and low error overestimation ratios (Figure S2a).
(a) $alpha=10.0$
(b) $alpha=1.0$
(c) $alpha=0.01$
(d) $alpha=0.001$
Figure S2: Module extraction runs with different regularization hyperparameter $\alpha$.

### v2

@@ A.1.2 Optimization of Locality Placement
We use Manhattan instead of Euclidean distances in the loss function to focus less on outliers and approximate centrality more equally across the samples (equivalent to the difference between minimizing the mean absolute and mean squared errors). The rationale for using the weighing scheme in Equation 1 is forcing the computationally cheap $\ell$-means algorithm to optimize our locality placement, i.e. minimize Equation 2. We want to weigh samples based on their centrality, since the $\ell$-means algorithm would otherwise minimize Euclidean distances rather than our objective. Furthermore, the weights in Equation 1 factor in the inverse average distances to offset the fact that $\ell$-means optimizes the sum of squared Euclidean distances between points, which means that lower density regions of the latent space might still have more probes (and localities), even when we weigh samples according to their centrality. Put together, we let the DIRECT algorithm optimize the weights of samples for the $\ell$-means algorithm, such that the localities produced by the $\ell$-means algorithm minimize our centrality-based loss function. Note that we use only the resulting centroids of the optimization as probes, while cluster assignments produced by the $\ell$-means algorithm are left unused. The $\ell$ localities are then simply the $n$-sized neighborhoods around each of the $\ell$ probes. Additionally, the DIRECT algorithm requires a range of values for the parameters in which the loss function is assumed to be convex. We fix this to $2D_{L},2D_{L}]$ for both $\alpha$ and $\beta$, where $D_{L}$ is the dimensionality of the latent space. We conjecture that, at least approximately (since $\ell$-means is stochastic, the loss function is convex in that interval and contains a good solution). Loosely, as mentioned above, we want the weighing scheme to be able to overpower the overrepresentation of low-density regions that might occur if we would simply weigh samples based on their centrality. We argue about it in terms of hyperball volume: the hyperball volume is proportional to the $D_{L}$-th power of its radius, which for a neighborhood of a sample we can approximate by the average distance to its neighbors, where $D_{L}$ is the dimensionality of the latent space. To battle overrepresentation, we then want to weigh samples with larges “volumes” (samples in low density regions) inversely to their volume, so as to effectively force $\ell$-means to give them less importance than they might otherwise have given the squared Euclidean distance optimization. In the two-dimensional case, the hyperball volume is simply the area of the circle, which is proportional to the square of the radius (the average distance of the $n$ nearest neighbors). To allow weighing in such a fashion, we allow $\alpha$ to be at least $2$, and we further multiply by $2$ to allow for exaggeration of this principle. Since we cannot be sure how this weighing will interact with centrality weighing, we allow both parameters to also be negative, and give them the same range, arbitrarily. In our limited experimentation, and with no ground truth to compare to, this seemed to yield well-placed localities results (Figures S12 and S28).
@@
The pseudocode, ommiting details about DIRECT, is shown in Algorithm S1. We use the scipy [^1] https://scipy.org/. implementation of the DIRECT algorithm, and the scikit-learn [^2] https://scikit-learn.org/. implementation of the $k$-means algorithm.
Algorithm S1 Locality Placement Algorithm.
require $\mathbf{X}\in\mathbb{R}^{E\times D_{L}}$, $\ell$
@@
A.1.3 Defining the Number of Localities
$\centering\ell=[E*o/n],\@add@centering$ (S1)
where [][\,] denotes rounding to the nearest integer. The hyperparameter $o$ determines the number of localities given a dataset and neighborhood size $n$. If all samples were of a similar centrality (e.g., if data are uniformly spread-out across some space), $o=x$ would mean that each sample would be represented by approximately $x$ localities. A larger $o$ will produce more overlap between localities and enable a representation that more closely approximates the original in-degree centralities of samples, but will be more expensive computationally and harder to manage in subsequent analysis. In Section 3 and Appendices B.5, we examine the stability of locality placement for UMAP embeddings of two different datasets in relation to hyperparameter $o$. Assuming stable (consistent) locality placement, we recommend using smaller values of $o$ in the range of $2$ to $20$, depending on the size of the dataset, so that the resulting number of localities $\ell$ is not above several thousand, but still captures the various correlation patterns and how they change across the latent space.
@@ A.2 Locality Representation
Correlations of each locality are calculated using the standard formulation of the Spearman correlation:
$r_{s}(\mathbf{X}_{i},\mathbf{X}_{j})=\frac{cov\left[\text{R}(\mathbf{X}_{i}),\text{R}(\mathbf{X}_{j})\right]}{\sigma_{\text{R}(\mathbf{X}_{i})},\sigma_{\text{R}(\mathbf{X}_{j})}},$ (S2)
where $X_{i}$ and $X_{j}$ are two original input features and R is a ranking function. Covariance covcov and standard deviations $\sigma$ are estimated from the samples within a locality using standard estimators. Tied ranks are averaged.
Importantly, if one of the variables has the same value in $75\%$ of the neighborhood, that correlation is set to 0 for that locality. This is done to prevent overestimation of correlations in cases where there is very little change. Notably, this has been the case for both the MNIST and KPMP datasets in our experiments, as the values of most pixels in MNIST are 0, and gene expression at the single-cell level in KPMP (and more generally) is very sparse. The value of $75\%$ was set arbitrarily, to allow for some leniency, while avoiding meaningless correlations based on lower numbers of samples.
We also note that other measures of association, and even univariate measures of variation, could be used depending on the intended use case. However, the different locality representation might require significant alternations to the subsequent module extraction.
@@ A.3.1 AMF loss function and hyperparameters
Different metrics of (dis)similarity. We decide to use cosine distance for the main term of our loss function. However, a different metric (or divergence) for assessing multivariate (dis)similarity might be equally suitable, e.g., the Kullback-Leibler divergence. We did not find strong argumentation about whether the semantics of one of these might be more appropriate for our task than the other. Work relating to semantics and efficiency of these metrics also suggests “triangular distance” as a more appropriate choice in some contexts (Connor, 2016), but is seldom used in contemporary literature.
Up-weighing overestimation and connection to quantile regression. The $\nu$ hyperparameter in Equation 6 determines how up-weighted overestimation errors will be in our loss. Weighing the errors is done similarly to how it is done in univariate quantile regression. The loss for univariate $\tau$-th quantile regression, also known as the pinball loss (Steinwart and Christmann, 2011), is defined as:
$\mathcal{L}_{\tau}(y,f(x))=\begin{cases}\tau\,|y-f(x)|&\text{if }y\geq f(x)\\ (1-\tau)\,|y-f(x)|&\text{if }y<f(x)\end{cases}.$ (S3)
Practically speaking, univariate quantile regression produces a weighing scheme between overestimation and underestimation errors: if $\tau>0.5$, we are less concerned with overestimating $y$ using $f(x)$, and vice versa. Using a $\tau=0.5$ is equal to doing median regression, which corresponds to a regression that minimizes the mean absolute error rather than using the more standard mean squared error. For more intuition, we refer to previous work (Koenker and Hallock, 2001; Steinwart and Christmann, 2011).
The idea for module extraction is to retain the semantics of similar, scale-invariant patterns we get with cosine distance or a similar metric, but allowing for reasoning about the reconstructed correlations in terms similar to quantile regression. Retaining the error weighing with a $\tau<0.5$ is therefore done to prevent averaging across samples and extract only shared, reoccurring correlations (Section 2.3). Quantile-regression-like reconstructions will allow us to say that correlations present in module-based reconstructions are actually present in the localities with some certainty, or, in terms of errors, that the value of a correlation will be overestimated by at most $\tau$ (approximately). We can quantify whether this is actually the case by looking at the overestimation errors (Figure 3b), which we calculate as the amount of overestimation errors divided by the total absolute error.
@@
$\centering\nu=\frac{1-\tau}{\tau}.\@add@centering$ (S4)
@@
Using $\nu$ for weighing overestimation errors when using the mean absolute errors would be equivalent to doing $\tau$-th quantile regression. We have set the hyperparameter $\nu=9$ in all of our experiments. If we look at the pinball equivalent $\tau$, this would correspond to $0.1$-th quantile regression. We cannot make direct parallels since we do not optimize the mean absolute error, but we can observe relative amounts of overestimation, which seem to more-or-less correspond to values we would expect if we would do quantile regression (Figure 3b).
@@
Extra: Presence fine-tuning. While we do not use it in our experiments, our implementation also includes what we call presence fine-tuning. After selecting modules, we can fix the module matrix $\mathbf{M}$ and optimize the presence matrix $\mathbf{P}$ with a pinball loss function (Equation S3), to guarantee the property that each locality (independent when $\mathbf{M}$ is fixed) has an overestimation ratio exactly equal to a chosen $\tau$. Notice that the modules we extract are still contingent on the overestimation weighing used in the main loss term (because modules and presences are optimized jointly), which means that using weighing in the main loss term is not equivalent to first extracting the modules without any constraints and then fine-tuning at the end.
@@
Scale regularization. The regularization term in Equation 6 is a simple mean absolute error term designed to retain scale of reconstructions. The idea behind scaling is simply to make reconstructions of the same scale as the actual locality correlations, as cosine distance is in itself scale invariant. While the cosine distance should remain relatively unaffected by this regularization, given a high enough hyperparameter $\gamma$, matching scale of the actual locality correlations and our reconstructions will allow us to observe the overestimation ratio of solutions (Section 3; Figure 3b). Furthermore, due to the scale invariance of cosine distance, it is also imaginable that modules avoid overestimation simply by retaining low correlations values across the board: the scaling will push optimization away from such solutions. It is crucial that the scale regularization contains no overestimation weighing, meaning that the ratio of overestimation is purely a result of the up-weighing in the first term of the loss function, as this ensures the retainment of the semantics of reoccurring patterns, rather than focusing on absolute values of our solutions. We set $\gamma=0.0001$ as the default value for scale regularization to have a low impact on the loss function. We also ended up manually setting this hyperparameter for the KPMP dataset to $\gamma=10.0$ in the experiments (see Section B.3.2), as we noticed $\gamma=0.001$ had no effect on optimization because of the low amounts of observed correlation.
@@
Optimization hyperparameters. The matrices $\mathbf{P}$ and $\mathbf{M}$ are optimized to minimize the loss function (Equation 6) using gradient descent with minibatches of $64$ samples (arbitrarily set), the ADAM optimizer (Kingma and Ba, 2017) with default hyperparameters, and clamping. Clamping is used to force any entries of matrices $\mathbf{P}$ and $\mathbf{M}$ to be within the $[0,1]$ interval after each step in training. Training is performed until convergence, which we define as no improvement in the loss function above $1\%$ in $100$ epochs. After convergence, we used the matrices from the iteration with the lowest loss as the final solutions as described in Section 3. We developed the code using the PyTorch library [^4] https://pytorch.org/..
@@ Appendix B Experiments
In this appendix we provide additional information about our experiments and show additional results and figures.
@@ B.1 Software and hardware
All our code was developed using Python 3.8 and its libraries [^5] Code included with supplementary material and will be made available online, conditional on acceptance.. We ran experiments on Linux compute nodes, comprising of a number of different CPUs and GPUs, details of which we omit to ensure anonymity. Locality placement and correlation calculations took several hours for both datasets. A single module extraction run would take about $5$ to $30$ minutes, increasing with the number of modules, $M$. We also tested all of our code on an Apple M3 Pro with 12 CPU cores and 18GB of RAM, which is what we also used to run Jupyter notebooks containing result analyses and plot generation.
@@ B.2.1 MNIST
The MNIST dataset was downloaded using the Scikit-learn Python library [^6] https://scikit-learn.org. library. No preprocessing was performed on the data. Prior to running LAVA, we created: (a) a 2-dimensional UMAP embedding of the dataset using the UMAP [^7] https://umap-learn.readthedocs.io. library, with default hyperparameters; (b) a 4-dimensional autoencoder embedding of the dataset using the denoising autoencoder from the Label-Free Explainability paper (Crabbé and Schaar, 2022).
@@
The KPMP single-cell kidney atlas was downloaded from the CELLxGENE website [^8] https://cellxgene.cziscience.com. Specific dataset version: https://datasets.cellxgene.cziscience.com/f5b6d620-76df-45c5-9524-e5631be0e44a.h5ad.. The dataset contains a UMAP embedding of the data preprocessed using the Seurat library [^9] https://satijalab.org/seurat/., which we used for our experiments. More information can be found at the KPMP data portal website [^10] https://atlas.kpmp.org/.. The UMAP of the data can also be explored on the CELLxGENE website [^11] https://cellxgene.cziscience.com/e/dea717d4-7bc0-4e46-950f-fd7e1cc8df7d.cxg/.. For module extraction, we removed mitochondiral, ribosomal, and mitochondrial ribosome genes. Then, we selected the top $1000$ most variable genes, based on the available gene expression data (data after normalization and log transformation).
@@ B.3.1 Size of locality, $n$
Data-driven approach (MNIST). A data-driven way of choosing the value $n$ can be looking at the Jaccard similarity of neighborhoods of varying sizes in the original space and the latent space. Even though methods such as UMAP have hyperparameters that represent the size of neighborhoods that will be optimized for, different neighborhood sizes might end up being optimized (depending on the choice of other hyperparameters and the dataset). For the UMAP of the MNIST dataset used in our experiments (Section 3), we found that with the default hyperparameter of $15$ nearest neighbors, the resulting embedding neighborhoods were most representative of the neighborhoods in the original feature space when $k\approx 3000$ (Figure S1), which was the value we used for LAVA’s hyperparameter in our experiments with MNIST. For simplicity, we also retained $k=3000$ for the LAVA experiments with the autoencoder-generated embeddings.
@@ B.3.2 Module extraction hyperparameters
For module extraction on both MNIST and KPMP, we use default hyperparameters, as defined in Appendix A.3, with one exception: the $\gamma$ regularization hyperparameter from Equation 6 in KPMP module extraction. In our runs, we noticed that using low values for $\gamma$ resulted in high overestimation error ratios for KPMP experiments. This pointed to the fact that regularization was mostly likely not affecting the loss function. We assumed this was due to the sparsity of gene expression, and subsequent low correlations in many localities. Running module extraction several times, using higher values of $\gamma$, confirmed this hypothesis (Figure S2). Increasing the strength of regularization seemed to have no or negative effect on the quality or stability of solutions (the main loss term and silhouette width, respectively). For our experiments, we ended up using $\gamma=10.0$, resulting in relatively low and stable error overestimation ratios (Figure S2a).
(a) $\gamma=10.0$
(b) $\gamma=1.0$
(c) $\gamma=0.01$
(d) $\gamma=0.001$
Figure S2: Module extraction runs with different regularization hyperparameter $\gamma$.